A bicycle chain of length 1.6 m and of mass 1 kg is lying on a horizontal floor. If g=10ms−2, the work done in lifting it with one end touching the floor and the other end 1.6 m above the floor is

Grade:12th pass

2 Answers

Vijay Kumar
7 years ago
mass of bicycle m = 1Kg
length h = 1.6 m
g = 10 m/s^2
Whenever the object like(ladder, stick, chain etc) are in contact with some other object or walls at the end points. its mass is divided into two equal parts.
work done W = mgh/2
therefore,
W = (1)(10)(1.6) / 2 = 8 J
dinesh sharma
7 years ago
If the chain is uniform and its one end is at ground and other is 1.6m above the ground.then its centre of mass is 0.8m above the ground.So change in potential energy = mgh = 1×10×0.8 = 8 joule = required work done